A projectile is thrown with speed ‘v’ at an angle  ‘x’ with upward vertical.It’s average velocity between the instants at which it crosses half of maximum height is?

Grade:12th pass

2 Answers

Vikas TU
7 years ago
Maximum height is given as:
H = u^2/2g
and half of its maximum height is: H/2 = > u^2/4g
Now u need to apply conservation of energy to find te veloocity at H/2.
U avg. = ucosthetha. i thinnk so.
Kalika Singh
5 years ago
H/2= V Sin2theta/4g
:- now you can find by placing values....
So, according to me answer is..Vsintheta , horizontal and in the plane of projections